OpenStudy (anonymous):

Solve the following solution? -2x+3y=3............ and x-6y=12

OpenStudy (anonymous):

solve for x in x-6y=12, take that answer and plug x into -2x+3y=3. The result will give you a y=... so then you can take that answer and plug it into x-6y=12

OpenStudy (anonymous):

thank so much

OpenStudy (anonymous):

solve first equation for x or y and then place its value in second equation like -2x+3y=3 solve it for x -2x=3-3y x=3-3y/-2 and put this value of x in second equation to get the value of y and then place the value of y in the first equation to get the value of x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you r welcum